What are the Z values of the limits of the area covering the middle 25% of the area under the normal curve?
Answer #1

We are to find k here such that:
P( -k < z < k) = 0.25

Due to symmetry about 0, we have here:
P( z < 0) = 0.5

Therefore, P(0 < z < k) = 0.25/2 = 0.125

Therefore P(z < k) = 0.5 + 0.125 = 0.625

From standard normal tables, we have:
P(Z < 0.3187) = 0.625

Therefore the z value limits here are given as:
(- 0.3187, 0.3187)